Solve and check the linear equation 12-6x=-54
Answer by macston(5194)

12-6x=-54 Subtract 12 from each side.
-6x=-66 Divide each side by -6.
x=11
CHECK:
12-6(11)=-54
21-66=-54
-54=-54